OpenStudy (anonymous):

35% of a number added to 39 is the number itself ,the number is ?

OpenStudy (lgbasallote):

it depends on the interpretation... one interpretation is .35x + 39 = x 39 = 0.65x x = 60 another interpretation is .35(x+39) = x .35x + 13.65 = x 13.65 = .65x x=21 dunno which you want :))